ABCD is a rectangle.If AB=7x-6 and CD=5x=30,find the value for x
Answer by jim_thompson5910(35256)

AB and CD are opposite sides. With any rectangle, opposite sides are congruent or equal in length.

So,

AB = CD

7x - 6 = 5x + 30 ... Note: I'm assuming you meant to say 5x + 30 instead of 5x = 30

Now we solve for x

7x - 6 = 5x + 30

7x = 5x + 30 + 6

7x - 5x = 30 + 6

2x = 36

x = 36/2

x = 18